OpenStudy (anonymous):

I am having a problem with method of substitution problem. 3x-y=6 7x-2y=16

OpenStudy (anonymous):

I am a little confused with the solution and I need help.

OpenStudy (roseflight):

Here's how I would do it: Change 3x-y=6 to 3x-6=y Then substitute the new y= for the y in the second equation like so: 7x-2(3x-6)=16 7x-6x+12=16 x+12=16 x=4 To find y you could then plug the x=4 into one of the first equations. Hope that helps :)

OpenStudy (anonymous):

Thank you so x=4 and y=6 Is this right?

OpenStudy (roseflight):

Yep, I think so.

OpenStudy (anonymous):

Thank you again..